# Isoperimetric inequalities and sharp upper bounds for Aharonov-Bohm eigenvalues on surfaces

### 13 Apr 2026 v1

@@ Appendix A Eigenvalues on the sphere
    $\{(\nu-n)(\nu-n+1)\}_{n\leq 0}$ $\text{ with multiplicity }|n|+1$
@@
In particular $\mu_{1}(\mathbb{S}^{2},A^{(\nu)}_{x_{0}})=\nu(\nu+1)$. If $\nu\in\mathbb{R}$, $\nu\notin\mathbb{Z}$, the formulas a) and b) hold with $\nu$ replaced by $\min_{m\in\mathbb{Z}}|m-\nu|$.
@@
Note that at the limit as $\nu\to 0$ we recover the spectrum of the Laplacian with multiplicities: the eigenvalues $(n-\nu)(n-\nu+1)$ go to $n(n+1)$ with multiplicity $n$ ($n\geq 1$) and the eigenvalues $(\nu-n)(\nu-n+1)$ go to $|n|(|n|+1)$ with multiplicity $|n|+1$ ($n\leq 0$). Altogether, the eigenvalues are given by $\ell(\ell+1)$ with multiplicity $2\ell+1$, $\ell=0,1,2,\cdots$.
@@
$\mu=\alpha(\alpha+1)$
Note that $P_{\alpha}^{-\alpha}(\cos(r))=c_{\alpha}\sin^{\alpha}(r)$, where $c_{\alpha}$ is a normalization constant.
Note also that there are two linearly independent solutions which are possibly singular at the endpoints $0,\pi$. From [13, §14.8, 14.8.1] we have
@@
It remains to consider the cases $\alpha=\pm|n-\nu|+k-1$, $k\in\mathbb{Z}$, $k\leq-1$. Note that if $\alpha=-|n-\nu|+k-1$ were a valid value, then a corresponding eigenvalue would be $\mu=\alpha(\alpha+1)=(-|n-\nu|+k-1)(-|n-\nu|+k)=(|n-\nu|+\ell)(|n-\nu|+\ell+1)$ with $\ell=-k\in\mathbb{Z}$, $\ell\geq 1$, which we already know to be in the spectrum. Hence we are left with the case $\alpha=|n-\nu|+k-1$, $k\in\mathbb{Z}$, $k\leq-1$. In this case, formula (A.4) is replaced by the following
@@
Step 2: compute multiplicities. To count multiplicities, we assume for simplicity that $\nu\in(0,\frac{1}{2}]$.
In this case, we see that the first eigenvalue is given by
@@
    $\{(\nu-n)(\nu-n+1)\}_{n\leq 0}$ $\text{ with multiplicity }{\frac{|n|+1}{2}}{\rm\ for\ }n\ {\rm odd\ and\ }{\frac{|n|}{2}}{\rm\ for\ }n\ {\rm even}$
@@
    $\{(\nu-n)(\nu-n+1)\}_{n\leq 0}$ $\text{ with multiplicity }{\frac{|n|+1}{2}}{\rm\ for\ }n\ {\rm odd\ and\ }{\frac{|n|+2}{2}}{\rm\ for\ }n\ {\rm even}$
@@
Note that the multiplicity of any Dirichlet/Neumann eigenvalue with $\nu=\frac{1}{2}$ is always an even number. For $\nu=0$ we find the multiplicities $n$ of the eigenvalue $n(n+1)$, $n\geq 0$, as Dirichlet eigenvalue of the Laplacian on $\mathbb{S}^{2}_{+}$, and the multiplicities $n+1$ of the eigenvalue $n(n+1)$, $n\geq 0$, as Neumann eigenvalue of the Laplacian on $\mathbb{S}^{2}_{+}$. Note that when the multiplicity is 0 we intend that the corresponding eigenvalue is not present in the spectrum.
@@
Let $u$ be the first eigenfunction on $\mathbb{S}^{2}$, associated with $\mu_{1}(\mathbb{S}^{2},A)$. From Proposition A.1, we know that it is radial and is given (up to constant multiples) by $u(r)=\sin^{\nu}(r)$ by the variational characterization of $\lambda_{1}(D_{R})$. Set $C_{R}\doteq D_{R}\setminus D_{2R-\pi}$ the spherical annulus obtained by removing from $D_{R}$ a the disk of radius $2R-\pi$ centered at $x_{0}$. Now, we have that $u\varphi\in H^{1}_{A}(D_{R})$ and from the min-max principle we have
@@
Everything is explicit, in fact $|d^{A}(u\varphi)|^{2}=|d(u\varphi)|^{2}+|A|^{2}u^{2}\varphi^{2}=u^{2}|\varphi|^{2}+\varphi^{2}|du|^{2}+\frac{\nu^{2}}{\sin^{2}(r)}u^{2}\varphi^{2}$. Standard computations show that
@@
On the other hand, let $u_{1}$ be an eigenfunction on $D_{R}$ associated with $\lambda_{1}(D_{R},A)$. Its extension by zero to $\mathbb{S}^{2}$ belongs to $H^{1}_{A}(\mathbb{S}^{2})$, hence, from the min-max principle for $\mu_{1}(\mathbb{S}^{2},A)$ we deduce that $\mu_{1}(\mathbb{S}^{2},A)\leq\lambda_{1}(D_{R},A)$. This concludes the proof. ∎
@@ Appendix B The spectrum of flat cylinders
Let $\mathcal{C}_{M}=\mathbb{S}^{1}\times[-M,M]$, with coordinates $(\theta,z)$, endowed with the flat metric $d\theta^{2}+dz^{2}$. Take the $1$-form $A=\nu d\theta$, which is closed and has flux $\nu$ around $\mathbb{S}^{1}$. If we think of $\mathcal{C}_{M}$ as a cylinder in $\mathbb{R}^{3}$, the potential $1$-form can be expressed in cartesian coordinates as $\frac{\nu}{x^{2}+y^{2}}(-ydx+xdy)$.
In order to compute $\mu_{1}(\mathcal{C})$, we write the magnetic Laplacian $\Delta_{A}$ in cylindrical coordinates. We obtain
@@
independently on $M$. The corresponding eigenfunctions are constant functions on $\mathcal{C}_{M}$.
The fact that the first eigenfunction of any flat annulus has constant modulus allows to prove an upper bound of isoperimetric type for any Riemannian cylinder, similarly to what happens to the magnetic Steklov problem in cylinders, see [15]. This is fundamnetal fact used in the proof of Theorem 2.6.
@@
Let $\mathcal{C}$ be a Riemannian annulus, namely, a differentiable surface diffeomorphic to $\mathbb{S}^{1}\times[0,1]$ endowed with a Riemannian metric. It is known that any Riemanniann annulus $\mathcal{C}$ is conformally equivalent to the flat cylinder $\mathcal{C}_{M}$ for a unique $M>0$, which is called the conformal modulus of $\mathcal{C}$. More explicitly, there exists a conformal diffeomorphism $\Phi:\mathcal{C}\to\mathcal{C}_{M}$. Consider on $\mathcal{C}_{M}$ the potential $1$-form $A=\nu d\theta$, which is closed and has flux $\nu$. Consider on $\mathcal{C}$ the potential $1$-form $\Phi^{\star}A$, which is again closed, and with flux $\nu$. It is unique up to exact forms. By gauge invariance, the magnetic spectrum of $\mathcal{C}$ depends only on $\nu$, see e.g., [8, §4] or [15] for more details. Let $u_{1}=1$. By the min-max principle we have
@@
As already mentioned, the spectrum of a Riemannian cylinder depends only on $\nu$, hence we may write $\mu_{1}(\mathcal{C},\nu)\doteq\mu_{1}(\mathcal{C},\Phi^{\star}A)$ and $\mu_{1}(\mathcal{C}_{M},\nu)\doteq\mu_{1}(\mathcal{C}_{M},A)$. If equality holds, then $u_{1}=1$ is an eigenfunction of $\Delta_{\Phi^{\star}A}$ on $\mathcal{C}$, which means that $|\Phi^{\star}A|^{2}=\nu^{2}$. This means that $\mathcal{C}$ supports a closed $1$-form of constant length. But since since the metric on $\mathcal{C}$ is conformal to the flat metric, we have that $|\Phi^{\star}A|^{2}=\rho^{2}|A|^{2}=\rho^{2}$ for some conformal factor $\rho$, and this implies that $\rho$ is constant. Hence $\mathcal{C}$ is homothetic to $\mathcal{C}_{M}$. ∎

### 29 Apr 2026 v2

@@ A.1. Spectrum of the punctured sphere
    $\{(\left|n\right|+\nu)(\left|n\right|+\nu+1)\}_{n\leq 0}$ $\text{ with multiplicity }|n|+1$
@@
  3. c)
    In particular, if $\nu\in(0,\frac{1}{2})$ one has μ_1(S^2,A^(ν)_x_0)=ν(ν+1) with eigenspace spanned by $u(r,\theta)=\sin^{\nu}(r)$. If $\nu=\frac{1}{2}$ the lowest eigenvalue is $\frac{3}{4}$, it is double, and the eigenspace is spanned by $\sin^{\frac{1}{2}}(r)$, $e^{i\theta}\sin^{\frac{1}{2}}(r)$.
  4. d)
    If $\nu\in\mathbb{R}$, $\nu\notin\mathbb{Z}$, formulae a), b) and c) hold with $\nu$ replaced by $\min_{m\in\mathbb{Z}}|m-\nu|$.
When $\nu\in(0,\frac{1}{2})$ the first few eigenvalues $\mu_{k}\doteq\mu_{k}(\mathbb{S}^{2},A^{(\nu)}_{x_{0}})$ are then given by
$\mu_{1}=\nu(\nu+1),\quad\mu_{2}=(1-\nu)(2-\nu)$
$\mu_{3}=\mu_{4}=(1+\nu)(2+\nu),\quad\mu_{5}=\mu_{6}=(2-\nu)(3-\nu)\quad\text{etc.}$
@@
Note that at the limit as $\nu\to 0$ we recover the spectrum of the Laplace-Beltrami operator with multiplicities: the eigenvalues $(n-\nu)(n-\nu+1)$ go to $n(n+1)$ with multiplicity $n$ ($n\geq 1$) and the eigenvalues $(\nu-n)(\nu-n+1)$ go to $|n|(|n|+1)$ with multiplicity $|n|+1$ ($n\leq 0$). Altogether, the eigenvalues are given by $\ell(\ell+1)$ with multiplicity $2\ell+1$, $\ell=0,1,2,\cdots$.
@@
When $\nu\in(0,\frac{1}{2})$, the first eigenvalue is attained at $n=0$, and equals
$\mu_{00}=\nu(\nu+1);$
it is simple, and the corresponding eigenspace is spanned by $\sin^{\nu}(r)$. When $\nu=\frac{1}{2}$ we have that the first eigenvalue is
$\mu_{00}=\mu_{10}=\frac{3}{4};$
it is double, and the corresponding eigenspace is spanned by $\sin^{\frac{1}{2}}(r)$ and $e^{i\theta}\sin^{\frac{1}{2}}(r)$.
@@
$\mu=\alpha(\alpha+1).$
Here $L^{b}_{a}$ denotes the Legendre function of degree $a$ and order $b$, see [13, §14]. Note that $P_{\alpha}^{-\alpha}(\cos(r))=c_{\alpha}\sin^{\alpha}(r)$, where $c_{\alpha}$ is a normalization constant.
Note also that there are two linearly independent solutions of the differential equation in (A.2) which are possibly singular at the endpoints $0,\pi$. From [13, §14.8, 14.8.1] we have
@@
It remains to consider the cases $\alpha=\pm|n-\nu|+k-1$, $k\in\mathbb{Z}$, $k\leq-1$. We discard the case $\alpha=-|n-\nu|+k-1$: in fact, a corresponding eigenvalue would be $\mu=\alpha(\alpha+1)=(-|n-\nu|+k-1)(-|n-\nu|+k)=(|n-\nu|+\ell)(|n-\nu|+\ell+1)$ with $\ell=-k\in\mathbb{Z}$, $\ell\geq 1$, which we already know to be in the spectrum. Hence we are left with the case $\alpha=|n-\nu|+k-1$, $k\in\mathbb{Z}$, $k\leq-1$. In this case, formula (A.4) is replaced by the following
@@
Step 2: compute multiplicities. To count multiplicities, we assume for simplicity that $\nu\in(0,\frac{1}{2}]$. We know that the first eigenvalue is given by
@@ A.2. Spectrum of the punctured hemisphere
    $\{(\left|n\right|+\nu)(\left|n\right|+\nu+1)\}_{n\leq 0}$ $\text{ with multiplicity }{\frac{|n|+1}{2}}{\rm\ for\ }n\ {\rm odd\ and\ }{\frac{|n|}{2}}{\rm\ for\ }n\ {\rm even}$
@@
    $\{(\left|n\right|+\nu)(\left|n\right|+\nu+1)\}_{n\leq 0}$ $\text{ with multiplicity }{\frac{|n|+1}{2}}{\rm\ for\ }n\ {\rm odd\ and\ }{\frac{|n|+2}{2}}{\rm\ for\ }n\ {\rm even}$
@@
Note that the multiplicity of any Dirichlet/Neumann eigenvalue with $\nu=\frac{1}{2}$ is always an even number. For $\nu=0$ we find the multiplicities $n$ of the eigenvalue $n(n+1)$, $n\geq 0$, as Dirichlet eigenvalue of the Laplacian on $\mathbb{S}^{2}_{+}$, and the multiplicities $n+1$ of the eigenvalue $n(n+1)$, $n\geq 0$, as Neumann eigenvalue of the Laplacian on $\mathbb{S}^{2}_{+}$. By our convention, if the multiplicity is 0 then the corresponding eigenvalue is not present in the spectrum.
A.3. Limit of Dirichlet eigenvalues
@@
Let $u$ be the first eigenfunction on $\mathbb{S}^{2}$, associated with $\mu_{1}(\mathbb{S}^{2},A)$. From Proposition A.1, we know that it is radial and is given (up to constant multiples) by $u(r)=\sin^{\nu}(r)$. Set $C_{R}\doteq D_{R}\setminus D_{2R-\pi}$ the spherical annulus obtained by removing from $D_{R}$ a the disk of radius $2R-\pi$ centered at $x_{0}$. Now, we have that $u\varphi\in H^{1}_{0,A}(D_{R})$ and from the min-max principle we have
@@
Everything is explicit, in fact $|d^{A}(u\varphi)|^{2}=|d(u\varphi)|^{2}+|A|^{2}u^{2}\varphi^{2}=u^{2}\varphi^{2}+\varphi^{2}|du|^{2}+\frac{\nu^{2}}{\sin^{2}(r)}u^{2}\varphi^{2}$. Standard computations show that
@@
On the other hand, let $u_{1}$ be an eigenfunction on $D_{R}$ associated with $\lambda_{1}(D_{R},A)$. Its extension by zero to $\mathbb{S}^{2}$ belongs to $H^{1}_{0,A}(\mathbb{S}^{2})$, hence, from the min-max principle for $\mu_{1}(\mathbb{S}^{2},A)$ we deduce that $\mu_{1}(\mathbb{S}^{2},A)\leq\lambda_{1}(D_{R},A)$. This concludes the proof. ∎
@@ Appendix B The spectrum of flat cylinders
Let $\mathcal{C}_{M}=\mathbb{S}^{1}\times[-M,M]$, with coordinates $(\theta,z)$, endowed with the flat metric $d\theta^{2}+dz^{2}$. Take the $1$-form $A=\nu d\theta$, which is closed, co-closed and has flux $\nu$ around any boundary component of $\mathcal{C}_{M}$. If we think of $\mathcal{C}_{M}$ as a cylinder in $\mathbb{R}^{3}$, the potential $1$-form can be expressed in cartesian coordinates as $\frac{\nu}{x^{2}+y^{2}}(-ydx+xdy)$.
In order to compute $\mu_{1}(\mathcal{C}_{M},A)$, we write the magnetic Laplacian $\Delta_{A}$ in cylindrical coordinates. We obtain
@@
independently on $M$. The corresponding eigenspace is spanned by the constant function $u_{1}=1$.
The fact that the first eigenfunction of any flat annulus has constant modulus will imply an upper bound of isoperimetric type on any Riemannian cylinder, similarly to what happens to the magnetic Steklov problem in cylinders, see [15]. This fundamental fact is used in the proof of Theorem 2.6.
@@
Let $\mathcal{C}$ be a Riemannian annulus, namely, a differentiable surface diffeomorphic to $\mathbb{S}^{1}\times[0,1]$ endowed with a Riemannian metric. It is known that any Riemanniann annulus $\mathcal{C}$ is conformally equivalent to the flat cylinder $\mathcal{C}_{M}$ for a unique $M>0$, which is called the conformal modulus of $\mathcal{C}$. In other words, there exists a conformal diffeomorphism $\Phi:\mathcal{C}\to\mathcal{C}_{M}$. Consider on $\mathcal{C}_{M}$ the potential $1$-form $A=\nu d\theta$, which is closed and has flux $\nu$. Consider on $\mathcal{C}$ the potential $1$-form $\Phi^{\star}A$, which is again closed, and with flux $\nu$. It is unique up to exact forms. By gauge invariance, the magnetic spectrum of $\mathcal{C}$ depends only on $\nu$, see e.g., [8, §4] or [15] for more details. Let $u_{1}=1$. By the min-max principle we have
@@
As already mentioned, the Aharonov-Bohm spectrum of a Riemannian cylinder depends only on the flux $\nu$, hence we may write $\mu_{1}(\mathcal{C},\nu)\doteq\mu_{1}(\mathcal{C},\Phi^{\star}A)$ and $\mu_{1}(\mathcal{C}_{M},\nu)\doteq\mu_{1}(\mathcal{C}_{M},A)$. If equality holds, then $u_{1}=1$ is an eigenfunction of $\Delta_{\Phi^{\star}A}$ on $\mathcal{C}$, which means that $|\Phi^{\star}A|^{2}=\nu^{2}$. This means that $\mathcal{C}$ supports a closed $1$-form of constant length. But since since the metric on $\mathcal{C}$ is conformal to the flat metric, we have that $|\Phi^{\star}A|^{2}=\rho^{2}|A|^{2}=\rho^{2}$ for some conformal factor $\rho$, and this implies that $\rho$ is constant. Hence $\mathcal{C}$ is homothetic to $\mathcal{C}_{M}$. ∎